OpenStudy (anonymous):

Please help me solve this multi step equation -7 + x = x/10 - 16

OpenStudy (anonymous):

I guess i would begin by removing the fraction. Multiply every term by (10/1). (10/1)(-7) + (10/1)x =(10/1)(x/10) -(10/1)(16) this becomes -70 +10x = x -160

OpenStudy (anonymous):

Then gather like terms -70 +(70) + 10x - x = x - (x) - 160 + (70) 9x = - 90 Divide both sides by 9 x = -10

OpenStudy (anonymous):

Sound good?

OpenStudy (anonymous):

That all looks correct!

OpenStudy (anonymous):

Thanks, Got it.